If 226 of 300 randomly selected newspaper subscribers indicated that they read the comic sections daily, estimate the probability that anyone subscriber selected randomly would not read the comics section daily


1
Expert's answer
2022-01-25T16:33:05-0500

"300-226=74" randomly selected newspaper subscribers indicated that they don't read the comic sections daily.

The probability that anyone subscriber selected randomly would not read the comics section daily is "\dfrac{74}{300}=\dfrac{37}{150}\approx0.2467."